OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 3, – 25), and x intercepts at x = -8 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 6x − ___ HELP!

jimthompson5910 (jim_thompson5910):

y = a(x-h)^2 + k is the basic vertex form where (h,k) is the vertex The given vertex is (-3,-25), so h = -3 and k = -25

jimthompson5910 (jim_thompson5910):

So y = a(x-h)^2 + k becomes y = a(x-(-3))^2 + (-25) y = a(x+3)^2 - 25

jimthompson5910 (jim_thompson5910):

Finally, a = 1 since the leading coefficient of y = x2 + 6x − ___ is 1

jimthompson5910 (jim_thompson5910):

so y = 1(x+3)^2 - 25 y = (x+3)^2 - 25

jimthompson5910 (jim_thompson5910):

Now expand and simplify

OpenStudy (anonymous):

How would I expand that? and what not.

jimthompson5910 (jim_thompson5910):

y = (x+3)^2 - 25 y = (x+3)(x+3) - 25 y = x(x+3)+3(x+3) - 25 Distribute that and combine like terms

OpenStudy (anonymous):

Would part of the answer be x^2+6x-16?

jimthompson5910 (jim_thompson5910):

yes it is, so the term that goes in the blank is 16

OpenStudy (anonymous):

Makes sense, thank you!

jimthompson5910 (jim_thompson5910):

yw